Suppose A is a 4 * 4 matrix such that det (A) = 16. Find the value of det[4(A^-1)^T ]


1
Expert's answer
2020-09-17T17:13:34-0400

Given that,


Determinant of a 4x4 matrix A, |A| = 16


By using the properties of determinants of matrices we can say that,

Determinant of any matrix = Determinant of its transpose

|XT| = |X|


Hence in our case we can say that,

|4(A-1)T| = |4(A-1)|


By using the property,

|k*X| = kn * |X| where k = constant & n = order of the square matrix, X


Hence in our case, n = 4

"\therefore" |4(A-1)| = 44 * |A-1|


But using the property,

"|X^{-1}| = \frac {1}{|X|}"


Substituting the values in above equation we get,

44 * |A-1| = 44 * "\frac {1}{16}" = 16


Hence,

If "|A|=16,\ \ \ \ |(A^{-1})^{T}|=16"